OpenStudy (anonymous):

Put the equation in the form ax2+bx+c=c. x2+6x+8=3

OpenStudy (saifoo.khan):

x^2 +6x + 5

OpenStudy (anonymous):

what about the equal sign?

OpenStudy (saifoo.khan):

x^2 +6x + 5 = 0

OpenStudy (anonymous):

x2+6x+5.5=5.5

OpenStudy (saifoo.khan):

y u wrote 5.5?

OpenStudy (anonymous):

because the question asks ax2+bx+c=c so both the digits before the equal sign and after the equal sign should be equal

OpenStudy (saifoo.khan):

whts the point of 5.5?